Question 5: (40 pts) Find the currents if(t) and i2(1) in the circuit shown below. 500 uF 12 10 cos 103: V 84mH
Answer #1

32 Jurc = Jx108x500u =-ja 21, 100 at W = 10 rady sec jwL=jx10240 =j4 32 + TOLO let vz be the voltage across 04.12 By applying

3.38 +j8.83 =vp[0-33 +056] Vg = 6.60 --j1867 views V3 = 7.05 |-15-369 se from circuit . in 1010*– 13 h : = 10 +jo- (6.60 -j1•